OpenStudy (anonymous):

How many grams of ethane gas (C2H6) are in a 12.7 liter sample at 1.6 atmospheres and 24°C?

OpenStudy (kainui):

\[\rm pressure \times volume = number ~of ~moles \times gas~constant ~\times temperature~in~Kelvin \]

OpenStudy (kainui):

convert atm to Pa and volume to m3 temperature to Kelvin

OpenStudy (kainui):

^ you'll be able to see that clearly if you know the gas constant.